If a coin is flipped and a die is rolled.what is the probability of getting heads and a 6?
The probability of flipping a heads is 1/2 and the probability of rolling a 6 is 1/6. By the laws of probability it would be logical to multiply them together, (1/2)(1/6) thus the answer being 1/12 with is roughly eight percent.
If the pack is well shuffled, the probability is 1/52.
What is the purpose of snowball sampling?
Snowball sampling can be used in a number of circumstances.
Sometimes the researcher may be able to identify only a small number of individuals in the target group but these people are likely to known others in the target group, and people that they know will also know more such people-and so on. By including those few initial people in her sample, and asking them to recruit others, the researcher may be able to achieve a useable sample size.
Then, sometimes, a researcher is actually more interested in the connections between people, or in other words, social networks. In this case, snowball sampling can be used to identify those connections.
Please see the link.
What is the probability of drawing 2 hearts and then a spade in that order from a deck of 52 cards?
In order to determine the probability of drawing 2 hearts and then a spade, in that order, from a deck of 52 cards, start by considering the first card.
The probability of drawing a heart is 1 in 4.
Since you have now reduced the number of hearts and the number of cards in the deck by one, the probability of drawing another heart is 4 in 17.
Since you have further reduced the number of cards by one, the probability of drawing a spade is 13 in 50.
Multiply these probabilities together, (1/4) (4/17) (13/50), and you get about 0.0153, or about 153 in 10000.
How many ways are there to select 10 balls from unlimited nunber of red white blue and green balls?
in those 10 balls we can have any combination...I am using here multinomial...
X=No. of red balls(0,1...10)
y=No. of green balls(0,1...10)
z=No. of white balls(0,1...10)
we need co-efficient of x^10 in (x^0+x^1+...+x^10)^3
that is (10+3-1)C3=12C3
Arrangement is also considered here
What is the probability of not tossing three heads with three fair coins?
The probability of something NOT happening is the complement of the probability of something happening. Since the probability that you DO have 3 heads is 1/8 (that is, 1/2 cubed), the complement is 1 - 1/8 = 7/8.
What is the probability of rolling a sum greater than 3?
90.47619%
Close but no cigar. I presume u r talking about rolling a pair of dice. There are 6 possible outcomes 4 each die, all equally likely, for a total of 36 equally likely possible outcomes. There is only 1 with a sum of 2, and 2 with a sum of 3, for a total of 3 outcomes with a sum of 3 or less. That's 3 out of 36, or 1/12. The other 11/12 represent a sum of more than 3, and that's 91.6666666...%
Probability of getting a perfect score in a three question true or false quiz?
If you answer randomly, 1 in 8.
If there is 5 true or false questions on a test how many possible out comes are there?
There are 2^5 outcomes or 32.
In how many ways can 5 speakers be seated in a row on a stage?
Imagine the 5 seats on the stage being filled in some order. The first seat can be occupied by any one of the 5 speakers. With each choice of 5 speakers, there are four who can go for the next seat. Then, wiyth each of the 5*4 choices, there are 3 speakers left to go for the next.
Continuing in this way, the answer is 5*4*3*2*1 = 5! = 120 ways.
In a standard deck of 52 cards, there are 13 clubs and 1 seven of hearts.
The probability of drawing a club or the seven of hearts, then, is 14 in 52, or 7 in 26.
It is 156/663 = 0.2353, approx.
A biased spinner has the numbers 1 2 3 4 what is the probability of getting a 4?
Without information about the bias, there can be no possible answer. You cannot even say that the probability of 4 is not 0.25 since suppose the spinner has the following probabilities:
Pr(1) = 0.1 Pr(2) = 0.4 Pr(3) = 0.25 Pr(4) = 0.25
is clearly biased - in favour of 2, but the probability of 4 is not affected by the bias.
What is the z score if 30 percent of the curve lies to the left of the z?
The z-score is 0.84. In the related link, look in the body of the table for .3 area (.2995 closest) and it yields the .84 z-score.
Two fair dice are rolled and the sum rolled is recorded find the probability that the sum is 4?
The two dice sum of 4:
1,3
2,2
3,1
= 3 in 36
= 1 in 12
What is the meaning of Sampling distribution of the test statistic?
Given any sample size there are many samples of that size that can be drawn from the population. In the population is N and the sample size in n, then there are NCn, but remember that the population can be infinite.
A test statistic is a value that is calculated from only the observations in a sample (no unknown parameters are estimated). The value of the test statistic will change from sample to sample. The sampling distribution of a test statistic is the probability distribution function for all the values that the test statistic can take across all possible samples.
How many different ways can the letters of the word CHORD be arranged?
5 letters in first position there can be any one of the five, in the second position there are four possibilities left, in third position one of three remaining letters etc. 5 * 4 * 3 * 2 *1 = 120.
If this is a homework assignment, please consider trying to answer it yourself first, otherwise the value of the reinforcement of the lesson offered by the assignment will be lost on you.
The range of probability for any event is between 0 and 1, inclusive.
If a die is rolled 3 times what's the chance of getting atleast one 6?
1/7776
50/50, or one chance in each set of three rolls. No matter how many times it's rolled, every roll is one in six. Just like flipping a coin. it's 50/50 heads or tails, even if you got 'heads' 20 times in a row the next flip will still be 50/50 chance of 'heads' again
What is the probability that a state does not begin with a vowel?
The answer depends on the state of which country!